A 27 kg chair initially at rest on a horizontal
floor requires a 176 N horizontal force to set
it in motion. Once the chair is in motion, a
148 N horizontal force keeps it moving at a
constant velocity.
The acceleration of gravity is 9.81 m/s
2
.
a) What is the coefficient of static friction
between the chair and the floor?
1
Expert's answer
2019-12-23T11:38:58-0500

Let the coefficient of static friction and the coefficient of kinetic friction is "\mu_s" and "\mu_k" .

N=mg

The condition to set in the motion, the applied force = friction force

i)

"F=\mu mg=\mu_s\times 27\times 9.81"


"\mu_s=\dfrac{176}{264.87}=0.66"

ii)

"\mu_k=\dfrac{F}{mg}=\dfrac{148}{27\times 9.81}=0.56"